A rigid vessel of volume 1m^3


1
Expert's answer
2021-07-29T14:25:01-0400

Question: A sealed rigid vessel has volume of 1 m3 and contains 2 kg of water at 100°C. The vessel is now heated. If a safety pressure valve is installed, at what pressure should the valve be set to have a maximum temperature of 200°C?


Answer:

Process: v = V/m = constant

State 1: v1 = 1/2 = 0.5 m3/kg from it's 2-phase

State 2: 200°C, 0.5 m3/kg

between 400 and 500 kPa so interpolate P ≅ "400 + 0.5-0.53422\/0.42492-0.53422 \u00d7 (500-400)" = 431.3 kPa